The human ear canal is about 2.8 cm long and
can be regarded as a tube open at one end and
closed at the eardrum.
What is the fundamental frequency around
which we would expect hearing to be most
sensitive? Assume the speed of sound in air
to be 340 m/s.
Answer in units of kHz.
1
Expert's answer
2020-02-21T10:06:25-0500
"f=\frac{v}{\lambda}=\frac{v}{4l}"

"f=\frac{340}{4(0.028)}=3036\ Hz"